The real risk-free rate is 3.71%, inflation is expected to be 3.21% this year, and the maturity risk premium is 0.7% while the liquidity premium and default risk premium is zero. IBM stock has a risk premium of 0.9%. What is the equilibrium rate of return on a 10-year Treasury bond?

Please round your answer to the fourth decimal. e.g. if your answer is 1.11%, you should input 0.0111.
Answer #1

In order to calculate the rate of return, nominal risk free rate should be known.

Nominal risk free rate is calculated below:

Nominal\, risk\, free\, rate=(1+real\, risk\: free\, rate)\times (1+inflation\, rate)-1Nominal\, risk\, free\, rate=(1+0.0371)\times (1+0.0321)-1

Nominal\, risk\, free\, rate=0.0704

The equilibrium rate of return is calculated below:

Rate\, of\, return=Maturity\, risk\, premium+Nominal\, risk\, free\, rate

Rate\, of\, return=0.007+0.0704

Rate\, of\, return=0.0774